OpenStudy (anonymous):

Given the points (2,1) and (3,4), answer the following. A) What is the slope of the line through these points? B) Write the equation in point slope form using the slope you found in part A and one of the ordered pairs given. It doesn’t matter which ordered pair you use. The point slope formula is (y – y1) = m(x – x1). C) Simplify your answer in part B and write the equation in slope-intercept form. Show all work.

OpenStudy (anonymous):

slope= (4-1)/(1-3)= -3/2 b. y-1=-2/3(x-2) c. y= -2/3x+ 7/3

OpenStudy (anonymous):

can u please jut help me answer it u can explain it to me later ok

OpenStudy (tennistar):

Do you have Geogebra?

OpenStudy (tennistar):

or any other graphing software?

OpenStudy (anonymous):

no :(( idkw ha u are tlaking about @ tennistar???

OpenStudy (tennistar):

geogebra is a software where you put in the points and it automatically shows it to u. Ok,  Do you know this?